OpenStudy (anonymous):

differentiate. fifth root of ln(X) ...i know its the product rule but im stuck on 1/5x + ?

OpenStudy (anonymous):

plezse shows steps

OpenStudy (kainui):

 d/dx(f(g(x)) = f'(g(x))*g'(x) right? 1/5x(ln(x))^(-4/5) = f'(g(x)) 1/x = g'(x) (1/5x(ln(x))^(-4/5)) * (1/x)